An initial amount of $300 accumulates to $381.25 after 6 years at a nominal rate of discount payable 4 times per year. The equivalent effective annual rate of interest is denoted i. Compute the accumulated value of $1000 invested for 30 months at a rate of simple interest of i per annum.
Answer #1

Güven that An initial amount of $300 accumbater to $ 381-25 after 6 years Accumulated value of $1900 invested for 30 months -dir-(.- 0.0396) 1 -0.960989 di: 0.639016 Then, i = 0.039016 : 1-0.239016 į =0.0406 Then Accamalled value of $1000 for 30 montCM 11 A = 1000 (14 010184481 293) = 1000(1+0.0461L5) = 1000 X11046115 JA = $1046.115 The Accumalaked walue of $ 1000 invested